OpenStudy (anonymous):

Magnitude of vector a is 2, magnitude of vector b is 3 and angle between them when placed tail to tail is 45 degrees. What is a.b ?

OpenStudy (anonymous):

tell how to calculate so i calculate others

OpenStudy (anonymous):

The dot product of two vectors a = [a1, a2, ... , an] and b = [b1, b2, ... , bn] is defined as: \[ \mathbf{a}\cdot \mathbf{b} = \sum_{i=1}^n a_ib_i = a_1b_1 + a_2b_2 + \cdots + a_nb_n\]

OpenStudy (anonymous):

and what about angle asad?

OpenStudy (anonymous):

\[ \mathbf{a} \cdot \mathbf{b}=\left\|\mathbf{a}\right\| \, \left\|\mathbf{b}\right\| \cos \theta \,\]